I believe that the Xend aux laser was turned off earlier today, for Steve's work swapping out the oplev. When I went down there, the red "off" LED was illuminated, and the LCD screen was showing something. I pushed the green "on" button, and I immediately got green.
Also, I saw that the 24Hz roll mode was very rung up on ETMX. I looked at the FM5 "bounce roll" filter, and it had some old values, 12Hz and 18Hz for the resonant gains. All other optics have the proper 16Hz and 24Hz frequencies. I copied the BS oplev bounce roll filter over to ETMX pit and yaw (both were wrong), and loaded them in. The mode is starting to ring down.